OpenStudy (anonymous):

Solve for x: -3 |x + 7| = -12

OpenStudy (campbell_st):

divide both sides of the equation by -3 and you get |x + 7| = 4 because you have an absolute value equation you have to solve 2 equations... x + 7 = 4 x + 7 = -4 hope it helps you can substitute you answers to check if they are correct